### v5

# Arena-Lite: Efficient and Reliable Large Language Model Evaluation via Tournament-Based Direct Comparisons

Abstract
As Large Language Models (LLMs) expand across domains, LLM judges have become essential for systems evaluation. Current benchmarks typically compare system outputs against baselines. This baseline-mediated approach, though convenient, yields lower reliability than direct comparison between systems. We propose Arena-Lite which integrates tournament structure on top of head-to-head comparison. The application of a tournament structure and direct comparison eliminates the need for baseline outputs, reduces the number of required comparisons, and allows higher reliability in system rankings. We conducted two experiments: (1) controlled stochastic modeling and (2) empirical validation with a real LLM judge. Those experiments collectively demonstrate that Arena-Lite consistently achieves higher reliability with fewer comparisons, even with smaller datasets or weaker judges. We release an easy-to-use web demonstration and code to foster adoption of Arena-Lite, streamlining model selection across research and industry communities. Arena-Lite demo and code are available on https://huggingface.co/spaces/NCSOFT/ArenaLite
Arena-Lite: Efficient and Reliable Large Language Model Evaluation via Tournament-Based Direct Comparisons
Seonil Son1††thanks: Current institution: RLWRLD Inc., Ju-Min Oh1††thanks: Current institution: Samsung AI Research, Heegon Jin1††thanks: Current institution: Global AI Platform, Cheolhun Jang1, Jeongbeom Jeong1, Kuntae Kim1††thanks: Current institution: Samsung Life Insurance 1NC AI Correspondence: seonil.calendar@gmail.com
1 Introduction
| Current Practice | ||||||
| Arena-Lite (ours) |
LLMs excel in diverse tasks, from chatbots to code generation, due to their powerful generative capabilities Ouyang et al. (2022); Roziere et al. (2023). As their versatility grows, accurately evaluating their performance becomes critical. To address this, benchmarks like MMLU and BigBench have emerged to assess LLM capabilities across various domains Hendrycks et al. (2020); Srivastava et al. (2023). Many of these benchmarks, such as those for arithmetic or code execution (e.g., GSM-Hard, HumanEval Gao et al. (2022); Chen et al. (2021)), use automated scoring to evaluate problem-solving skills. However, their focus is not on quality of generated content or limited to the cases where the generated contents are automatically evaluated (e.g. programming), which are mostly not the case for variety of generation tasks. The Chatbot Arena, a leading platform for reliable human evaluation of LLMs, has set a standard by collecting extensive human annotations Chiang et al. (2024). Yet, its resource-intensive approach has prompted efforts to replicate its rankings using LLM judges as a cost-effective alternative Li et al. (2024, 2023). These methods, however, rely on baseline-mediated comparisons—comparing LLM outputs to a leading proprietary model’s outputs—which sacrifice reliability.
Current benchmarks relying on baseline often rank LLMs by their win rate against baseline responses from an leading proprietary models. This approach has two advantages: it scales linearly with the number of LLMs and provides a consistent quality standard. However, we argue that comparing LLMs directly against each other is inherently more reliable than using baseline outputs, which can introduce noise coming from weak transitivity Xu et al. (2025) of human preferences on LLM responses. To address this, we propose Arena-Lite, a novel evaluation framework that uses direct, head-to-head comparisons organized in a tournament structure. By eliminating the need for baseline outputs, Arena-Lite reduces the number of comparisons required while achieving stronger alignment with human-established rankings, such as those from Chatbot Arena.
Arena-Lite conducts single-elimination tournaments among participating LLMs for each prompt. From the match results, we can compute Bradley-Terry preference ratings for the final ranking Bradley and Terry (1952). This results in a single scalar per model that captures relative performance between any model counterpart, enabling accurate and efficient ranking. We validate Arena-Lite through two experiments. The first experiment, stochastic modeling of LLM competition (Section 4.2) demonstrates that tournament-based direct comparison method outperforms baseline-mediated method under various conditions, including different numbers of LLM participants, dataset rows used, and judge accuracies. Second, our empirical experiment (Section 4.3) shows that Arena-Lite achieves higher correlation with Chatbot Arena’s rankings than standard approaches using baseline outputs (Table 1) attested toward number of LLM as judges. These results collectively highlight Arena-Lite’s ability to deliver reliable rankings with fewer comparisons, even with smaller datasets or weaker judges over various generation tasks.
Our contributions are threefold:
1.
We introduce Arena-Lite, a tournament-based framework for direct LLM comparisons, offering greater reliability than baseline-mediated approaches.
2.
We rigorously demonstrate, through both comprehensive modeling and empirical experiments, that Arena-Lite achieves more accurate LLM rankings while requiring fewer comparisons than prevalent practices, particularly those relying on common baseline model outputs.
3.
We open-source a functional demo and the complete code for Arena-Lite (https://huggingface.co/spaces/NCSOFT/ArenaLite), enabling researchers and industry practitioners to easily host and utilize our framework for streamlined LLM evaluation.
2 Preliminaries: Quantifying Generative Performance
Quantifying the generative capabilities of LLMs is an inherently challenging task. The evaluation is complicated by the stochasticity of model outputs and the subjectivity of human judgments. To approximate real-world performance, a common methodology involves assessing model outputs across a diverse range of prompts. In this context, two metrics are widely utilized: the win rate, which measures the frequency of preference for a model’s output over a baseline, and the Bradley-Terry model, which is employed to infer a latent skill rating for each model based on pairwise comparisons.
2.1 Measuring Win rate over baseline outputs
Benchmarks like AlpacaEval and Arena-Hard-Auto assess LLM response quality by comparing it to baseline responses from proprietary model Li et al. (2023, 2024). An LLM judge evaluates whether the candidate LLM’s response outperforms the baseline for a given prompt. The win rate—the proportion of prompts where the LLM’s response is preferred—serves as a measure of its generative ability. While this approach is straightforward and scalable, it introduces noise coming from mediated comparisons.
2.2 Bradley-Terry Model Preference for LLM Rating
The Bradley-Terry (BT) model Bradley and Terry (1952) is widely used to infer baseline-mediated rankings of LLMs from pairwise comparisons. Chatbot Arena adopts the BT model rather than the classical Elo system Elo and Sloan (1978), but both Elo and BT models are useful for expecting probability of match outcome based on a score difference, though they differ in update rules and statistical assumptions.
In the BT model, each LLM is assigned a latent score representing its procificency. Given LLMs and with scores and , respectively, the probability that LLM is preferred over LLM is modeled as:
This formulation closely resembles the Elo win-probability function, reinforcing the intuitive connection between the two.
Chatbot Arena uses this BT-based formulation to rank LLMs by aggregating human preferences collected through pairwise matchups Chiang et al. (2024). Users are shown responses from two anonymized models to the same prompt and asked to select which response they prefer. The accumulated judgments are then used to fit BT scores, producing a leaderboard that reflects relative model performance.
While this approach requires a substantial number of human evaluations to ensure reliability, it captures nuanced quality differences between models more effectively than purely automatic benchmarks. Arena-Lite, introduced in the next section, builds on the same BT modeling framework but seeks to reduce the number of required comparisons by using tournament-structured match-making.
3 Arena-Lite
To address the high annotation cost of Chatbot Arena while preserving evaluation reliability, we propose Arena-Lite. Arena-Lite introduces a tournament-based approach for efficient and reliable LLM evaluation using a single-elimination structure. Unlike baseline-mediated evaluations that compare model outputs to a baseline, Arena-Lite directly compares outputs from different models through head-to-head matchups for each prompt in benchmark datasets. Repeated tournaments across the dataset produce consistent leaderboards reflecting models’ fundamental performance.
We first discuss limitations of baseline-mediated evaluations (Section 3.1). Next, we describe how Arena-Lite conducts tournaments to generate ratings (Section 3.2, Algorithm 1). Finally, we highlight similarities between the single-elimination structure and merge sort, explaining why aggregated tournaments yield reliable LLM rankings (Section 3.3).
3.1 Comparing to Baseline outputs is not Always Helpful
Although baseline outputs are a standard way to evaluate and rank LLMs, they introduce potential failure modes. Beyond the fact that a single baseline output might not capture every dimension of appropriate answers, relying solely on a baseline output can lead to unreliable rankings of LLMs.
Consider an ideal scenario with a judge capable of perfectly distinguishing the quality of any two outputs. If we choose to compare LLM responses directly to rank them using BT preference (Equation 1), all head-to-head comparisons are utilized. In contrast, baseline-mediated evaluation for differentiating LLMs can exhibit failure modes, as shown in Equation 2.
When the baseline output () for a prompt () successfully disambiguates the pair of LLM responses and (as in the first and second cases), comparison to the baseline is effective for benchmarking. Otherwise, these comparisons do not help differentiate LLM performance. Consequently, the baseline-mediated approach provides less information for ranking when multiple responses are either both correct or both incorrect relative to the baseline.
3.2 Tournaments of LLMs over multiple prompts to preference ratings
Figure 1 and Algorithm 1 illustrate how Arena-Lite benchmarks LLMs via a tournament approach. Here, denotes the number of prompts in the benchmark dataset. Running Arena-Lite hosts tournaments among participant LLMs for every prompt in the dataset.
The use of tournament structures for LLM benchmarking offers both benefits and challenges. A major advantage of a single-elimination tournament is efficiency. As shown in Table 1, the number of matches scales linearly with the number of participants and even lower compared to using baseline outputs. However, single elimination tournament only identifies a champion, leaving the relative ordering of other participants unclear.
To retain tournament’s efficiency while obtaining a fine-grained ranking, we propose aggregating tournament results over multiple prompts with randomized initial match-ups for each prompt. Performing multiple tournaments with random initialization offers several benefits:
1.
It resolves ties among non-champion participants from previous tournaments.
2.
It mitigates the impact of unfavorable match-ups in any single tournament.
3.
Aggregating match results allows for precise win rate estimation via BT preference, resulting in a well-aligned overall ranking.
4.
More matches are allocated to high-performing participants while ensuring every participant is evaluated at least once per prompt (Table 1).
In Section 3.3, we further explain how aggregating multiple tournaments could yield reliable ranking of LLMs. We also provide further analysis on number of comparisons performed over tournaments of Arena-Lite comparing to merge sort, offering a comprehensive view of the method’s efficiency and effectiveness.
3.3 Why Aggregating Multiple Tournaments Yields Reliable Rankings
A key challenge in LLM evaluation is to derive a reliable ranking from a feasible number of pairwise comparisons. Our tournament-based approach, Arena-Lite, is designed to efficiently sample these comparisons. In a single tournament with n models, each model participates in a minimum of one match and a maximum of matches. When conducted over a benchmark with distinct prompts, the total number of evaluations per model is bounded within the interval as presented in Table 1. However, the efficiency of this tournament structure raises a critical question: how does this limited sampling produce a statistically reliable global ranking?
To achieve reliable rankings of LLMs, our approach aggregates match outcomes from multiple tournaments, lesser than a full grid comparisons, but effectively approximating the complete set of pairwise comparisons required for merge sort. We outline the rationale in four key points:
Merge Sort Baseline A single-elimination tournament mirrors the merging steps of merge sort, which requires comparisons with no duplicate match-ups to rank models. However, a single tournament omits many comparisons, covering only the minimal match-ups needed to determine a winner.
Recovering Comparisons via Aggregation By aggregating tournaments over diverse prompts, we can recover missed pairwise match-ups that should have occurred. Assuming match outcomes are prompt-independent (as BT model assumes), matches across prompts are considered equivalent. With prompts (typically hundreds to thousands) and models (tens), only random initial match-ups totaling . This exceeds the possible combination, ensuring broad coverage.
Sufficiency of Comparisons The aggregated match-ups not only cover the necessary comparisons but also surpass the requirement of merge sort. Moreover, each unique model pair competes in average to matches111These estimates are computed by number of matches a model undergoes which is , devided by number of possible opponents for a model, across the benchmark, a frequency mostly sufficient for accurate win rate estimation.
Refinement for Reliability The remaining matches, totaling , further refine the ranking by enhancing win rate estimates, especially among top-performing models, reducing noise and ensuring robustness akin to Arena-Lite’s sampling strategy.
In summary, aggregating multiple tournaments reconstructs the full set of comparisons needed for a merge sort-like ranking while providing enough repeated match-ups to ensure accurate win rate estimations. This dual mechanism yields reliable and robust LLM rankings across the benchmark.
4 Experiments
We conducted two experiments to evaluate Arena-Lite against baseline-mediated benchmarking. The first experiment (Section 4.2) utilized a stochastic model to simulate LLM competitions, comparing Arena-Lite’s tournament-based direct comparison with baseline-mediated evaluation. This controlled setup allowed us to test Arena-Lite’s design principles, such as the effectiveness of direct versus mediated comparison (Section 3.1) and tournament-based sampling (3.3), while isolating variables and minimizing noise, such as LLM judge biases Park et al. (2024). The second experiment (Section 4.3) validates Arena-Lite empirically using various LLMs as judges and public benchmark data. We tested models including gpt-4o, gpt-4o-mini, Claude3.5, Qwen2.5, Llama3.1, and Gemma2 to assess Arena-Lite’s effectiveness against standard benchmarking practices. Together, these experiments demonstrate the superior reliability and efficiency of Arena-Lite’s tournament approach. Section 4.1 outlines shared experimental settings, followed by detailed descriptions of each experiment in subsequent subsections.
4.1 Chatbot Arena Leaderboard as Ground-Truth Rankings
We benchmark Arena-Lite and baseline-mediated evaluation against rankings from the Chatbot Arena leaderboard, widely recognized for its reliability due to extensive human preference annotations. With a large volume of votes across diverse prompts, these rankings provide a robust ground truth for model comparisons.
4.2 Experiment 1: Controlled Stochastic Modeling of LLM Competitions
We suggest a simple stochastic model based on the Bradley-Terry (BT) framework to compare Arena-Lite’s approach with baseline-mediated evaluation. This experiments simulates prompt-agnostic LLM competitions which follows the Bradley-Terry model’s presumption, with outcomes determined by a judge following Equation 3. The judge’s decision is based on the BT preference difference () between models and , and the judge’s accuracy ():
With the model of judge above (Equation 3), we simulate both Arena-Lite’s tournament-based approach and baseline-mediated approaches according to the following initial conditions and procedures.
Initial conditions:
Ground-Truth BT Preference: We extracted BT preferences from the English category of Chatbot Arena (as of June 23), derived from approximately 60% of user-submitted judgments. These preferences serve as both the initial model parameters and the ground-truth rankings for evaluation.
Judge Accuracy (): We varied judge accuracy from 0.6 to 0.9.
Number of LLMs () and Dataset Size (): We adjusted the number of participating LLMs and benchmark dataset sizes to assess the robustness of both approaches in data-poor and data-rich settings.
Simulation Procedure:
1.
Select participant LLMs and their BT preferences.
2.
Compute expected win rates () using Equation 3.
3.
Sample match outcomes based on (Equation 3), determined by the BT preference rating gap () and judge accuracy ().
4.
Repeat for the specified number of test prompts ().
5.
Compute scores:
Baseline-mediated: Win rate against a reference model (gpt-4-1106-preview, rating 1233).
Arena-Lite: BT preference from all tournament match outcomes.
6.
Rank models based on scores.
7.
Calculate Spearman correlation between simulated and ground-truth rankings.
We conducted 50 trials per configuration to account for stochasticity in initial tournament brackets and judging process.
4.3 Experiment 2: Empirical Validation of Arena-Lite with real LLM Judge
To empirically validate our proposal, we evaluated the reliability of both Arena-Lite and baseline-mediated approach over the top 19 models from the Chatbot Arena leaderboards. This experiment employs actual prompt inputs and LLM outputs, distinguishing it from the earlier simulation study.
4.3.1 Dataset: Test Prompts and LLM Responses Used
Testing the benchmarking approaches requires: (1) test prompts and (2) the corresponding responses from LLMs. For the benchmark dataset, we selected Arena-Hard-Auto Li et al. (2024). The prompts in Arena-Hard-Auto were carefully curated from Chatbot Arena user queries. This dataset consists of 500 prompts—two instances for each of 250 subtopics. Although AlpacaEval Li et al. (2023), which comprises 800 prompt-reference pairs, could serve as a viable testbed, we opted for Arena-Hard-Auto because its design aligns more closely with Chatbot Arena. Arena-Hard-Auto uses responses from gpt-4-0314 as the baseline outputs. For ranking, we utilized the reserved outputs of the top 20 (=19 + baseline) models from the Arena-Hard-Auto Browser.222Extracted from the 2024 Jul 6 commit (fd42026).
4.3.2 Participant LLMs
For ranking, we selected 19 LLMs from the top of the ChatBot Arena leaderboard in the hard prompts category, as these models most closely align with Arena-Hard-Auto.
4.3.3 LLM Judges
We used several aligned LLMs as judges for testing both benchmarking approaches.
LLMs of our choice are gpt-4o family of models OpenAI et al. (2024), Claude3.5, and a selection of open-weight models: Qwen2.5 Qwen et al. (2025), Llama3.1 Grattafiori et al. (2024), and Gemma2 Team et al. (2024).
For pairwise comparisons of responses, we employed the judging prompt suggested in LLMBar Zeng et al. (2024) (See Appendix A.8.2).
The same judge prompt was applied consistently across both the tournament and baseline-mediated approaches.
To mitigate position bias Wu and Aji (2023), the order of model responses was alternated during evaluation.
Further details on the LLM-as-a-judge configuration are provided in Appendix A.8.
The two experimental settings are summarized as follows:
Experiment 1 (Modeling Experiment): This experiment uses the ground truth BT preference of the models to initialize the simulation. We vary control parameters for the benchmarking approaches—including the judge’s accuracy (), the number of test prompts used (), and the number of participant LLMs ()—to determine which benchmarking approach more accurately reproduces the participants’ ranking. For each configuration, we conduct 50 trials of experiments.
Experiment 2 (Empirical Validation): This experiment assesses the two benchmarking approaches using empirical runs with various LLM judges. We select the top 19 LLMs from Chatbot Arena and used their reserved outputs on Arena-Hard-Auto test prompts. For both the tournament and baseline-mediated approaches, we employ the Spearman correlation coefficient to measure how well the results align with the ground truth leaderboard rankings. In our empirical study, we conduct 500 trials for each experimental setting.
5 Results and Discussion
We assess the reliability and robustness of Arena-Lite as a means for LLM benchmarking, comparing it against the current baseline-mediated approach. Our results from both simulation study and empirical runs indicate that the tournament approach of Arena-Lite yields rankings that align more closely with the ground-truth Chatbot Arena leaderboards. We present our findings using whisker plots and tables in the following sections.
5.1 Experiment 1: Modeling Experiment Results
Figure 2 illustrates noticeable differences in Spearman correlation, indicating that the tournament approach is more reliable than the baseline-mediated method. The consistent performance gap across various conditions—namely, the number of participants, the number of test prompts, and judge accuracy (, , and )—demonstrates the robustness of the tournament approach. Although the simulation simplifies real-world complexity, a similar performance gap was observed in the empirical findings (Experiment 2, Figure 3). This consistency suggests that the robust performance of Arena-Lite is not coincidental or limited to a specific empirical setting of ours.
5.2 Experiment 2: Empirical Validation Results
As hinted in the previous section, the empirical results in Figure 3 show that Arena-Lite consistently outperforms the baseline-mediated approach. Although the performance gaps are less pronounced than in the simulation, the same trend persists. In Table 2, we report the median values for Arena-Lite and the baseline-mediated approach using the gpt-4o family of judges while varying the number of test prompts (). These results consistently demonstrate that Arena-Lite outperforms the baseline-mediated method. Note that Arena-Lite shows similar or superior reliability even in extreme data-poor benchmark condition ().
Table 3 presents the outcomes when using other LLMs as judges, with a fixed number of prompts (). The results for Claude3.5-sonnet, Llama3.1-8b, and Qwen2.5-7b follow a similar trend. However, smaller models (Gemma2-2b and Qwen2.5-0.5b) appears to be less reliable as an LLM judge. Hence, we recommend using evaluation-specialized judge LLMs or, at least, generative judge models with around 7B parameters regardless of using Arena-Lite or considering baseline-mediated approach.
| Spearman corr. () | = 50 | 100 | 250 | 475 | 500 |
| baseline-mediated (4o) | 0.895 | 0.935 | 0.963 | 0.966 | 0.964 |
| Arena-Lite (4o) | 0.905 | 0.940 | 0.960 | 0.970 | 0.970 |
| baseline-mediated (4o-mini) | 0.895 | 0.908 | 0.917 | 0.916 | 0.912 |
| Arena-Lite (4o-mini) | 0.901 | 0.919 | 0.931 | 0.933 | 0.933 |
| = 500 |
| baseline-mediated | 0.924 | 0.820 | 0.756 | 0.089 | 0.592 | ||||||||||
| Arena-Lite | 0.930 | 0.850 | 0.811 | -0.124 | 0.552 |
5.3 Incorporating a New LLM into an Existing Leaderboard
While our main focus has been on ranking multiple LLMs at once, it is also useful to consider the common scenario of adding a single new model to an existing leaderboard, which is also frequent use-case. We explored two approaches: (1) a binary search-like placement method, and (2) using the top-performing model response as a baseline. Our findings indicate that the later approach is more reliable (Table 5, Appendix). Further details and discussions are provided in Appendix A.6.
6 Related Works
6.1 LLM-as-a-Judge for Systems Ranking
Utilizing LLM-as-a-Judge as a building block for systems ranking has become a common practice in the LLM benchmarking community. Several studies have investigated how LLM judges compare to human evaluators, examining their similarities and differences Park et al. (2024), as well as how these differences impact system rankings (e.g., JuStRank Gera et al. (2024), Gao et al. (2025)). Our research extends these approaches by proposing a method that orchestrates LLM-as-a-Judge through a well-established tournament structure to derive rankings among systems.
6.2 Efficient and Reliable Evaluation
There is a growing body of research focused on optimizing the number of evaluations while maintaining reliability when using LLM-as-a-Judge for system ranking. Perlitz et al. proposed a metric called DIoR to quantify the relationship between computational costs and system ranking reliability. UniCBE Yuan et al. (2025) introduced a method to analyze the relationship between reliability and the number of judge evaluations based on uncertainty. BenchBench Perlitz et al. (2024b) systematically analyzed consistency across benchmarks and provided a package to facilitate this analysis. tinyBenchmarks Maia Polo et al. (2024) explored strategies to minimize the number of evaluations across various established benchmarks. Arena-Lite relates to these studies in that it leverages the properties of tournament structures and direct comparisons to achieve more reliable results with fewer judge evaluations.
7 Conclusion
We introduced Arena-Lite, an efficient and reliable framework for evaluating Large Language Models (LLMs) through tournament-based direct comparisons. By eliminating the need for baseline outputs and adopting head-to-head comparison, Arena-Lite achieves higher reliability in system rankings with reduced number of comparisons. Our experiments, encompassing controlled stochastic modeling and empirical validation with various LLM judges, confirm that Arena-Lite consistently outperforms standard baseline-mediated evaluation methods, even with smaller datasets or weaker judges. The release of an accessible web demonstration and code supports the adoption of Arena-Lite to help streamlining model development cycle across research and industry. Future work will extend Arena-Lite’s application to diverse domains, including multi-modal LLM evaluation involving visual or audio inputs and outputs.
Limitations
While we conducted extensive testing to assess the robustness of Arena-Lite tournaments—including 50 and 500 trials for Experiment 1 and Experiment 2, respectively—some inherent sources of randomness remain, such as variation due to initial match bracket assignments. The randomness in bracket assignment is added for adopting tournament structure of Arena-Lite and may influence outcome stability. Future work could explore more informative or adaptive matchmaking strategies that improve ranking fidelity beyond what is achievable with single-elimination formats, potentially within the same or even fewer number of matches.
Appendix A Appendix
A.1 Arena-Lite Web Demo
We provide screenshots of Arena-Lite web demo here (link to Huggingface Space). You could try or locally host Arena-Lite according to its documentation. It is quite easy to use and do not require much of resources to host.
Arena-Lite provides the benchmark result (Figure 4) with helpful visualization interface that enables walking through the matches and tournaments one by one (Figure 5) and match statistics between LLMs (Figure 6). We also provide visualization that helps examining potential bias of LLM Judge being used (Figure 7).
A.1.1 Starter Prompt Set
We provide several judge prompts that we have used for specific target tasks. Some of those are for quite specialized tasks, and some might work for evaluating general instruction following. You could customize your own judge prompt based on your evaluation needs according to the documentation. The list of the judge prompts we provide is as follows, and one could see the detailed prompts at yaml files here.
1.
llmbar prompt (Figure A.8.2) and conciser version of the prompt, llmbar_brief. Those are for evaluating instruction following.
2.
translation_pair prompt for selecting translation models trained on game-specialized parallel corpora,
3.
rag_pair_kr prompt for evaluating knowledge groundedness of korean RAG models over chatting scenario,
4.
translation_fortunecookie prompt which was crafted for evaluating translation models specialized for translating fortune-tellings, and
5.
post_edit prompt for evaluating conversation revision based on given persona of a speakers given in a instruction, which is quite specialized use case.
A.2 Full table for Experiment 2
Here is the extended results of Experiment 2 (Section 4.3) presented in Table 3. Aligned LLMs smaller than 7B parameters struggles to work as a proper Judge. Otherwise, Arena-Lite method excels over common practice of using outputs from proprietary as baselines.
| Dataset size | method | claude 3.5 | llama3.1 | qwen2.5 | qwen2.5 | gemma2 |
| sonnet | 8b-it | 7b-it | 0.5b-it | 2b-it | ||
| 50 | baseline-mediated | .896 | .656 | .492 | .010 | .064 |
| Arena-Lite (ours) | .897 | .715 | .544 | -0.051 | -0.088 | |
| 100 | baseline-mediated | .912 | .732 | .596 | .002 | .079 |
| Arena-Lite (ours) | .918 | .780 | .656 | -0.068 | -0.090 | |
| 250 | baseline-mediated | .924 | .801 | .700 | .045 | .560 |
| Arena-Lite (ours) | .929 | .830 | .760 | -0.131 | .551 | |
| 475 | baseline-mediated | .924 | .819 | .708 | .083 | .112 |
| Arena-Lite (ours) | .930 | .845 | .810 | -0.131 | -0.009 | |
| 500 | baseline-mediated | .924 | .820 | .756 | .089 | .592 |
| Arena-Lite (ours) | .930 | .850 | .811 | -0.124 | .551 |
A.3 Machine Requirements for Experiments
Except the part we inferenced open-weight models such as Llama, Qwen and Gemma, our experiments are mostly do not require GPU usage. Inference are done on one A100 GPU, but T4 would be enough for reproducing our experiments. Otherwise, our experiments require querying API and post-processing those with CPU. Experiments could be run on personal desktops. The lowest specification of the machine we deployed had i5-8400 CPU, 16 GiB RAM.
A.4 Assuring Statistical Significance of the Results within Budget for proprietary models
To ensure a statistically significant number of trials for each experiment while staying within budget, we utilize OpenAI’s Batch API to prepare full-grid match outcomes (i.e., all-play-all matches for every prompt) in a cache file, allowing us to reuse these outcomes. Each empirical experiment consists of 500 trials per setting, with results represented using whisker plots or summary statistics such as median values. When experimenting with a subset of the Arena-Hard-Auto benchmark (), we sample a stratified subset of the benchmark dataset for each new trial.
A.5 BT preference from Arena-Lite compared to Human Annotations
Figure 8 shows the BT preference computed out of Arena-Lite. For judge, we used gpt-4o. As mentioned in the caption, the BT preference are bootstrapped median value from 500 trials. 95% confidence intervals also plotted as an error bar, which look negligible in scale compared to observed values. Matches are performed over Arena-Hard-Auto benchmark dataset (500 prompts).
A.6 Binary search vs. Win rate over baseline
A.6.1 Binary Search
We tried binary search placement of a newly added LLM to the leaderboard without baseline output in Table 6. Details of how we implemented binary search are attached in Algorithm 2, Appendix. It turns out that binary search based on already built leaderboard ranks is not as reliable compared to utilizing the best model’s outputs as a baseline. Therefore, when adding a newcomer LLM to pre-existent leaderboard, we could utilize the already submitted responses as a baseline from the 1st placed LLM.
| () | gt=1-6 | 7-13 | 14-19 (20) | total avg. |
| binary search (4o) | 0.92 | 1.84 | 2.13 | 1.72 |
| comp. to 1st (4o) | 1.98 | 1.55 | 1.57 | 1.39 |
| binary search (4o-mini) | 1.27 | 1.82 | 1.21 | 1.5 |
| comp. to 1st (4o-mini) | 1.00 | 1.43 | 1.43 | 1.37 |
A.6.2 Comparing to the most Performant Model so far: Converting Ratings Table back to Win Rates
Assuming we preserved a set of match results and model outputs from the last benchmarking, we could benefit from those to perform insertion. One could pick an appropriate anchor LLM as a baseline in a leaderboard to estimate the skill of a newcomer. Using previous matches from the tournaments that built the leaderboard could be used for estimating win rates over the baseline. This is the same as converting the preference ratings table into a win rate leaderboard. Since the leaderboard is not built with full-grid matches but with tournaments, there would be some missing matches against the baseline regardless we have picked. There are two ways to estimate the win rate over the baseline model. We could just count the matches given are enough in amount, or we could also convert BT preference back to to use it directly for scoring for the model ranks in the leaderboard. Reminding that BT preference rating is for expecting a likely outcome of the match, this should work. After this win rate of the newcomer model could be directly compared for enlisting.
| () | gt=1 | 2 | 3 | 4 | 5 | 6 | avg. |
| binary search | 0.09 | 1.24 | 1.75 | 1.55 | 1.26 | 1.10 | 0.92 |
| (4o) | (.04/-.03) | (.14/-.14) | (.09/-.09) | (.07/-.06) | (.08/-.08) | (.10/-.09) | |
| anchored | 0.00 | 1.01 | 1.95 | 2.00 | 0.96 | 0.30 | 1.98 |
| (4o) | (0.00/0.00) | (0.01/-0.01) | (0.02/-0.02) | (0.00/0.00) | (0.02/-0.02) | (0.04/-0.04) | |
| binary search | 0.52 | 0.85 | 0.59 | 2.03 | 1.20 | 2.45 | 1.27 |
| (4o-mini) | (.09/-.07) | (.12/-.11) | (.10/-.09) | (.02/-.02) | (.05/-.05) | (.07/-.06) | |
| anchored | 0.00 | 0.00 | 1.00 | 2.00 | 2.00 | 1.00 | 1.00 |
| (4o-mini) | (0.00/0.00) | (0.00/0.00) | (0.00/0.00) | (0.00/0.00) | (0.00/0.00) | (0.00/0.00) |
| 7 | 8 | 9 | 10 | 11 | 12 | 13 | avg. |
| 1.31 | 1.27 | 2.22 | 1.74 | 2.27 | 2.23 | 1.86 | 1.84 |
| (.10/-.10) | (.11/-.11) | (.14/-.12) | (.09/-.09) | (.12/-.11) | (.12/-.12) | (.07/-.07) | |
| 0.30 | 3.68 | 1.09 | 1.03 | 2.97 | 0.78 | 1.00 | 1.55 |
| (0.04/-0.04) | (0.04/-0.04) | (0.03/-0.03) | (0.02/-0.01) | (0.02/-0.02) | (0.05/-0.05) | (0.00/0.00) | |
| 0.69 | 0.85 | 3.89 | 1.95 | 2.10 | 2.37 | 0.88 | 1.82 |
| (.07/-.06) | (.09/-.09) | (.12/-.11) | (.06/-.05) | (.03/-.03) | (.10/-.11) | (.12/-.11) | |
| 0.51 | 0.52 | 3.50 | 1.00 | 1.00 | 3.00 | 0.50 | 1.43 |
| (0.49/-0.51) | (0.48/-0.52) | (0.49/-0.51) | (0.00/0.00) | (0.00/0.00) | (0.00/0.00) | (0.50/-0.50) |
| 14 | 15 | 16 | 17 | 18 | 19 | 20 | avg. |
| 1.40 | 3.07 | 0.80 | 1.47 | 5.00 | 0.96 | - | 2.13 |
| (.04/-.05) | (.11/-.11) | (.08/-.09) | (.05/-.04) | (.11/-.11) | (.08/-.09) | ||
| 2.00 | 2.00 | 1.00 | 1.21 | 3.00 | 0.21 | - | 1.57 |
| (0.00/0.00) | (0.00/0.00) | (0.00/0.00) | (0.03/-0.04) | (0.00/0.00) | (0.04/-0.03) | ||
| 1.45 | 4.20 | 0.19 | 0.08 | 1.09 | 1.08 | 0.40 | 1.21 |
| (.07/-.08) | (.17/-.17) | (.07/-.06) | (.03/-.02) | (.05/-.05) | (.05/-.05) | (.07/-.07) | |
| 1.00 | 2.00 | 2.00 | 1.00 | 1.00 | 3.00 | 0.00 | 1.43 |
| (0.00/0.00) | (0.00/0.00) | (0.00/0.00) | (0.00/0.00) | (0.00/0.00) | (0.00/0.00) | (0.00/0.00) |
A.7 Separability In terms of Confidence Interval
To see how well the two benchmarking approach (anchored comparison and tournament approach) separates LLMs in adjacent ranks, we provide scatter plot of Elo rating and win rate paired with error bar (95% confidence interval). We present the both results of using gpt-4o (Figure 9) and gpt-4o-mini (Figure 9) as a judge. Inside the each plot, inseparables indicates the cases where any pair of datapoint co-includes each other within their range of error bars, and overlap means a certain datapoint is within some other’s range of error, when it is one-sided.
A.8 Judge configuration
A.8.1 Evaluation Prompt
A.8.2 Decoding Parameters
We did not configure decoding parameters of judge LLMs (gpt-4o[-mini]), which its temperature defaults to 1. The only parameter we have adjusted is maximum number of tokens to be generated, which for our prompt is less than 6 (i.e. The output of our prompt is (a) or (b)). To avoid position bias, we alternated the position of the responses from a certain model across the benchmark prompt.
PROMPTS = [ # metrics.txt from LLMBar
"role": "system",
"content": "You are a helpful assistant in evaluating the quality of the outputs for a given instruction. Your goal is to select the best output for the given instruction.",
"role": "user",
"content": """Select the Output (a) or Output (b) that is better for the given instruction. The two outputs are generated by two different AI chatbots respectively.
Here are some rules of the evaluation:
(1) You should prioritize evaluating whether the output honestly/precisely/closely executes the instruction, then consider its helpfulness, accuracy, level of detail, harmlessness, etc.
(2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction.
(3) You should avoid any potential bias and your judgment should be as objective as possible. For example, the order in which the outputs were presented should NOT affect your judgment, as Output (a) and Output (b) are **equally likely** to be the better.
Do NOT provide any explanation for your choice.
Do NOT say both / neither are good.
You should answer using ONLY "Output (a)" or "Output (b)". Do NOT output any other words.
# Instruction:
instruction
# Output (a):
response_a
# Output (b):
response_b
# Questions about Outputs:
Here are at most three questions about the outputs, which are presented from most important to least important. You can do the evaluation based on thinking about all the questions.
- Does the output well satisfy the intent of the user request?
- If applicable, is the output well-grounded in the given context information?
- Does the output itself satisfy the requirements of good writing in terms of:
1) Coherence
2) Logicality
3) Plausibility
4) Interestingness
# Which is better, Output (a) or Output (b)? Your response should be either "Output (a)" or "Output (b)":""",
] # prompt ends here
LLMBar prompt of our use. We used metric variant suggested in original LLMBar paper. More preset prompts are in our Arena-Lite Demo and source (https://huggingface.co/spaces/NCSOFT/ArenaLite)